100g of ice at 0 degree Celsius is mixed with 10g of steam at 100 degree Celsius. Find the final composition
 

Grade:9

2 Answers

Arun
4 years ago
Latent heat of ice =m×l 100×80=8000Latent heat of steam =m×l =10×540 L.H of ice is greater than steam Specific heat of steam = mct =10×1×100=10005400+1000=6400 for steamCompositionQ=mlm=q/lm=6400/80m=80100-80=20g of ice 10+80=90g water
Vikas TU
4 years ago
Dear student 
Kindly follow the below link to get ans of your question. 
Hope this helps 
Good Luck